# On local Galois deformation rings: generalised tori

### v1

@@ 1. Introduction
Let $G$ be a smooth affine group scheme over $\mathcal{O}$, such that its neutral component $G^{0}$ is torus and the component group $G/G^{0}$ is a finite étale group scheme over $\mathcal{O}$. We call such group schemes generalised tori. We do not make any assumptions regarding $p$ and the component group of $G$.
@@
There is a finite extension $L^{\prime}$ of $L$ with the ring of integers $\mathcal{O}^{\prime}$ and a continuous representation $\rho:\Gamma_{F}\rightarrow G(\mathcal{O}^{\prime})$ lifting $\overline{\rho}$.
@@
If $G\cong G^{0}\rtimes(G/G^{0})$ then a lift in Theorem 1.1 can be constructed over $\mathcal{O}$ using the Teichmüller lift to define a section of $G^{0}(\mathcal{O})\twoheadrightarrow G^{0}(k)$, which then induces a section $\sigma:G(k)\rightarrow G(\mathcal{O})$. Then $\sigma\circ\overline{\rho}$ is the required lift. In particular, in this case Theorem 1.1 holds if we replace $\Gamma_{F}$ with any profinite group. However, it seems non-trivial to prove 1.1 in general and our argument uses that the Euler–Poincaré characteristic formula holds for $\Gamma_{F}$.
@@
Assume that Theorem 1.1 holds with $\mathcal{O}=\mathcal{O}^{\prime}$ then
@@
The above theorems are used in an essential way in a companion paper [11], where we study deformations of Galois representations with values in generalised reductive $\mathcal{O}$-group schemes $G$, which means that the neutral component $G^{0}$ is reductive and the component group $G/G^{0}$ is finite étale. If we let $G^{\prime}$ be the derived subgroup scheme of $G^{0}$, then $G^{0}/G^{\prime}$ is a torus and $G/G^{\prime}$ is a generalised torus as considered in this paper. If $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ is a continuous representation then we show in [11, Proposition 4.20] that its deformation ring $R^{\square}_{\overline{\rho}}$ can be presented over the deformation ring of $\varphi\circ\overline{\rho}:\Gamma_{F}\rightarrow(G/G^{\prime})(k)$, where $\varphi:G\rightarrow G/G^{\prime}$ is the quotient map. This allows us to split up the arguments into “torus part”, carried out in this paper, and “semisimple part” carried out in [11].
@@ 1.1. Arithmetic setting
Let us consider the general case now. The first issue, alluded to in 1.2, is that it is not clear anymore that such a lift exists. However, let us postpone the explanation how we deal with this problem until Section 1.3 and let us give ourselves a lift $\rho_{0}:\Gamma_{F}\rightarrow G(\mathcal{O})$ of $\overline{\rho}$. Then the map $\rho\mapsto[\gamma\mapsto\rho(\gamma)\rho_{0}(\gamma)^{-1}]$ considered above induces a bijection between $D^{\square}_{\overline{\rho}}(A)$ and the set $\widehat{Z}^{1}(A)$ of continuous $1$-cocycles $\Phi:\Gamma_{F}\rightarrow\operatorname{Hom}(M,1+\mathfrak{m}_{A})$. Further, one may interpret $\widehat{Z}^{1}$ as $D^{\square}_{\overline{\rho}}$ for the group $G=\mathfrak{D}(M)\rtimes\underline{\Delta}$ and $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ given by $\overline{\rho}(\gamma)=(1,\pi(\gamma))$, where $\pi:\Gamma_{F}\rightarrow\Delta$ is the quotient map. So that $\widehat{Z}^{1}$ is the analogue of $D^{\square}_{\mathbf{1}}$ and we have to find the ring pro-representing the functor $\widehat{Z}^{1}$.
@@
Our solution is motivated by Langlands paper [10] on his correspondence for tori, where he shows that the universal coefficient theorem induces an isomorphism
@@
corresponding to the fundamental class $[u_{E/F}]\in H^{2}(\Delta,E^{\times})$. Since every continuous representation $\Gamma_{F}\rightarrow G(A)$ factors through the profinite completion of $W_{E/F}$, this motivates us to study the functor
@@ 1.2. Abstract setting
is representable by the group algebra $\mathbb{Z}[H_{1}(\Gamma_{1},M)]$.
@@
To prove 1.7 we verify that the action of $\Delta$ on $(\Gamma_{2}^{\mathrm{ab}}\oplus I_{\Delta})\otimes M$ induced by its action on the functor in (6) is isomorphic to $\mathcal{E}\otimes M$. This calculation is carried out in Sections 4. The last part of Theorem 1.7 follows by observing that (1) remains an isomorphism if $\mathbb{C}^{\times}$ is replaced by any divisible group and for any $A$ there is a faithfully flat map $A\rightarrow B$ such that $B^{\times}$ is divisible.
@@
$\operatorname{Rep}^{\Gamma_{1}}_{G}\sslash G^{0}\overset{\cong}{\longrightarrow}\mathrm{PC}^{\Gamma_{1}}_{G}$
@@
Let us assume that $\operatorname{Rep}^{\Gamma_{1}}_{G}(\mathcal{O})$ is non-empty, $G^{0}$ is split over $\mathcal{O}$ and $G/G^{0}$ is a constant group scheme. Let $\rho_{0}:\Gamma_{1}\rightarrow G(\mathcal{O})$ be a representation and let $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ be the subfunctor of $\operatorname{Rep}^{\Gamma_{1}}_{G}$ consisting of representations $\rho$ such that $\Pi\circ\rho=\Pi\circ\rho_{0}$, where $\Pi:G\rightarrow G/G^{0}$. One may similarly define $\mathrm{PC}^{\Gamma_{1}}_{G,\pi}$. We show that the map
@@ 1.3. Producing a lift
Let us go back to the problem of exhibiting a lift of $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ to the characteristic zero as claimed in Theorem 1.1. Using $\overline{\rho}$ instead of $\rho_{0}$, we may identify the restriction of $D^{\square}_{\overline{\rho}}$ to $\mathfrak{A}_{k}$ with the restriction of $\widehat{Z}^{1}$ to $\mathfrak{A}_{k}$. This allows us to conclude that
@@
$\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}(\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M\cdot([F:\mathbb{Q}_{p}]+1),$
@@ 1.5. Notation
We denote by $\operatorname{Set}$, $\operatorname{Ab}$, $\operatorname{CRing}$ the categories of sets, abelian groups and commutative rings, respectively.
If $X$ is a scheme the we denote the ring of its global sections by $\mathcal{O}(X)$. If $\Gamma$ is a group then we denote by $\mathcal{O}[\Gamma]$ its group algebra over a ring $\mathcal{O}$.
@@ 1.6. Acknowledgements
VP would like to thank Timo Richarz for a helpful discussion concerning the Langlands correspondence for tori. The authors would like to thank Toby Gee and James Newton for their comments on a draft version of this paper.
@@ 2. Generalised tori
It follows from [5, Proposition 3.1.3], that for a generalised torus $G$ the quotient $G/G^{0}$ is étale. Sean Cotner has shown in [6] that if $S$ is locally noetherian then for a smooth affine $S$-group scheme $G$ the identity component $G^{0}$ has reductive geometric fibers and $G/G^{0}$ is finite étale if and only if $G$ is geometrically reductive in the sense of [1, Definition 9.1.1]. In particular, this holds over $S=\operatorname{Spec}\mathcal{O}$.
@@
If $M$ is an abelian group then we denote by $\mathfrak{D}(M):=\operatorname{Spec}(\mathcal{O}[M])$ the diagonalizable group scheme, where the Hopf algebra structure on $\mathcal{O}[M]$ is given by addition in $M$ and inversion. For any $\mathcal{O}$-algebra $A$, we have
@@ 3. Shapiro’s lemma
Let $\varphi:\Gamma_{2}\rightarrow V$ be a $1$-cocycle, and let $f_{\varphi}:\Gamma_{2}\rightarrow\mathcal{F}_{V}$ be a $1$-cocycle corresponding to $\varphi$, via the canonical isomorphism $\mathcal{F}_{V}\cong V$ and let ${\Phi_{\varphi}:\Gamma_{1}\rightarrow\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}{V}}$ be a function given by
@@
The map $\varphi\mapsto\Phi_{\varphi}$ depends on the choice of the coset representatives $\bar{c}$ and so it is not canonical. However once these representatives have been fixed it is immediate from the formulas that the map is functorial in $V$. So that if $\alpha:V\rightarrow W$ is a $\Gamma_{2}$-equivariant homomorphism of abelian groups equipped with $\Gamma_{2}$-action and $\psi\in Z^{1}(\Gamma_{2},W)$ is the image of $\varphi$ under the map $Z^{1}(\Gamma_{2},V)\rightarrow Z^{1}(\Gamma_{2},W)$ induced by $\alpha$ then $\Phi_{\varphi}$ maps to $\Phi_{\psi}\in Z^{1}(\Gamma_{1},\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}{W})$.
@@ 4. The space of $1$-cocycles
Let $V$ be an abelian group equipped with a trivial $\Gamma_{2}$-action. We may identify $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V$ with the space of functions $f:\Delta\rightarrow V$ and $[g\cdot f](c)=f(cg)$ for all $g\in\Gamma_{1}$ and $c\in\Delta$. We also have an action of $\Delta$ on $\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}V$, which commutes with the action of $\Gamma_{1}$ and is given by $[d\cdot f](c)=f(d^{-1}c)$ for all $d\in\Delta$. This induces an action of $\Delta$ on $Z^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V)$, which is given explicitly by
@@
We want to understand the action of $\Delta$ on $B^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V)\oplus\operatorname{Hom}(\Gamma_{2},V)$ via this isomorphism. It follows from (8) that $[[d\ast\Phi](g)](cd)=[\Phi(g)](c)$. Hence
$[d\ast\Phi_{\varphi}](g)=\sum_{c\in\Delta}\varphi(\overline{c}g\overline{cg}^{-1})\mathbf{1}_{cd}.$
@@
$[[d\ast\Phi_{\varphi}](k)](1)=\sum_{c}\varphi(\overline{c}k\overline{c}^{-1})\mathbf{1}_{cd}(1)=\varphi(\overline{d^{-1}}k\overline{d^{-1}}^{-1})=\varphi(\overline{d}^{-1}k\overline{d})=[d\cdot\varphi](k).$
We note that $\varphi((\overline{d}h)^{-1}k(\overline{d}h))=\varphi(\overline{d}^{-1}k\overline{d})$ for all $h,k\in\Gamma_{2}$, which justifies the last two equalities. Thus the difference $d\ast\Phi_{\varphi}-\Phi_{d\cdot\varphi}$ is a $1$-coboundary and we want to find a function $f:\Delta\rightarrow V$ depending on $d$ such that for all $g\in\Gamma_{1}$ we have
@@
Since $V$ is abelian we have $\operatorname{Hom}(\Gamma_{2},V)=\operatorname{Hom}(\Gamma_{2}^{\mathrm{ab}},V)$, where $\Gamma_{2}^{\mathrm{ab}}$ is the maximal abelian quotient of $\Gamma_{2}$. For $d,c\in\Delta$ we define $\kappa(d,c)$ to be the image of $\overline{d}\overline{c}\overline{dc}^{-1}$ in $\Gamma_{2}^{\mathrm{ab}}$. Then $\kappa:\Delta\times\Delta\rightarrow\Gamma_{2}^{\mathrm{ab}}$ is the $2$-cocycle corresponding to the extension
@@
see [3, Chapter XIV, §4, Remark]. The extension class of (17) gives an element in $\operatorname{Ext}^{1}_{\mathbb{Z}[\Delta]}(I_{\Delta},\Gamma_{2}^{\mathrm{ab}})$. One can show, that the image of this class in $H^{2}(\Delta,\Gamma_{2}^{\mathrm{ab}})$ is equal to the class of $\kappa$.
@@
where the action of $\Gamma_{1}$ on the last term is induced its action on $\mathbb{Z}[\Delta]\otimes M$, which is given by $g\cdot(c\otimes m)=cg^{-1}\otimes m$. The diagonal action of $\Delta$ on $\mathbb{Z}[\Delta]\otimes M$ is given by $d\cdot(c\otimes m)=dc\otimes dm$, commutes with the action of $\Gamma_{1}$ and induces an action of $\Delta$ on $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V)$, which commutes with the action of $\Gamma_{1}$. If $M=\mathbb{Z}$ with the trivial $\Delta$-action then the construction recovers the action of $\Delta$ on $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V$ considered earlier.
@@
$(\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V))^{\Delta}$ as $\Gamma_{1}$-representation is naturally isomorphic to $\operatorname{Hom}(M,V)$ with $\Gamma_{1}$ acting on it via $\Delta$.
@@
Let $b\in B^{1}(\Gamma_{1},\operatorname{Hom}(M,V))$ be a boundary $b(g)=(g-1)\alpha$. Its image in $B^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V))$ is the boundary $b^{\prime}$ given by $b^{\prime}(g)=(g-1)\vartheta(\alpha)$. The constant function $f^{\prime}(c)=\alpha$ for all $c\in\Delta$ is $\Gamma_{1}$-invariant thus
@@
If $V$ is any abelian group then for all $\Gamma_{1}$-modules $N$ we have a canonical map
(21) $H^{1}(\Gamma_{1},\operatorname{Hom}(N,V))\rightarrow\operatorname{Hom}(H_{1}(\Gamma_{1},N),V),$
@@
Lemmas 4.6 and 4.8 give us the commutative diagram above. Moreover, if $V$ is divisible then the top row is exact, the last arrow in the top row is surjective and the first vertical arrow is an isomorphism. We deduce that for all divisible $V$ the maps $q:(\mathcal{E}\otimes M)_{\Delta}\twoheadrightarrow I_{\Delta}M$ and $\varphi_{\mathrm{nat}}:H_{1}(\Gamma,M)\rightarrow(\mathcal{E}\otimes M)_{\Delta}$ induce an exact sequence
@@
(25) $\mathbb{Z}[(\mathcal{E}\otimes M)_{\Delta}]\cong\mathbb{Z}[H_{1}(\Gamma,M)][t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}],$
@@
The right hand side is an fpqc sheaf, since it is representable [14, Tag 023Q]. So it suffices to find for every ring $A$ and every homomorphism ${f:\mathbb{Z}[H_{1}(\Gamma_{1},M)]\to A}$ an fpqc cover of $A$ and a descent datum in the left hand side, which maps to a descent datum in the right hand side, which descends to $f$. By [16, Lemma 4.1.1] for every ring $A$ there is a faithfully flat map $A\to B$, such that $B^{\times}$ is divisible. By 4.7 the map

### v2

@@ 1. Introduction
Let $G$ be a smooth affine group scheme over $\mathcal{O}$, such that its neutral component $G^{0}$ is a torus and the component group $G/G^{0}$ is a finite étale group scheme over $\mathcal{O}$. We call such group schemes generalised tori. We do not make any assumptions regarding $p$ and the component group of $G$.
@@
There is a finite extension $L^{\prime}$ of $L$ with ring of integers $\mathcal{O}^{\prime}$ and a continuous representation $\rho:\Gamma_{F}\rightarrow G(\mathcal{O}^{\prime})$ lifting $\overline{\rho}$.
@@
If $G\cong G^{0}\rtimes(G/G^{0})$ then a lift as in Theorem 1.1 can be constructed over $\mathcal{O}$ using the Teichmüller lift to define a section of $G^{0}(\mathcal{O})\twoheadrightarrow G^{0}(k)$, which then induces a section $\sigma:G(k)\rightarrow G(\mathcal{O})$. Then $\sigma\circ\overline{\rho}$ is the required lift. In particular, in this case Theorem 1.1 holds if we replace $\Gamma_{F}$ with any profinite group. However, it seems non-trivial to prove 1.1 in general and our argument uses that the Euler–Poincaré characteristic formula holds for $\Gamma_{F}$.
@@
Assume that Theorem 1.1 holds with $\mathcal{O}=\mathcal{O}^{\prime}$. Then
@@
The above theorems are used in an essential way in a companion paper [12], where we study deformations of Galois representations with values in generalised reductive $\mathcal{O}$-group schemes $G$, which means that the neutral component $G^{0}$ is reductive and the component group $G/G^{0}$ is finite étale. If we let $G^{\prime}$ be the derived subgroup scheme of $G^{0}$, then $G^{0}/G^{\prime}$ is a torus and $G/G^{\prime}$ is a generalised torus as considered in this paper. If $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ is a continuous representation then we show in [12, Proposition 4.20] that its deformation ring $R^{\square}_{\overline{\rho}}$ can be presented over the deformation ring of $\varphi\circ\overline{\rho}:\Gamma_{F}\rightarrow(G/G^{\prime})(k)$, where $\varphi:G\rightarrow G/G^{\prime}$ is the quotient map. This allows us to split up the arguments into “torus part”, carried out in this paper, and “semisimple part” carried out in [12].
@@ 1.1. Arithmetic setting
Let us consider the general case now. The first issue, alluded to in 1.2, is that it is not clear anymore that such a lift exists. However, let us postpone the explanation how we deal with this problem until Section 1.3 and let us give ourselves a lift $\rho_{0}:\Gamma_{F}\rightarrow G(\mathcal{O})$ of $\overline{\rho}$. Then the map $\rho\mapsto[\gamma\mapsto\rho(\gamma)\rho_{0}(\gamma)^{-1}]$ considered above induces a bijection between $D^{\square}_{\overline{\rho}}(A)$ and the set $\widehat{Z}^{1}(A)$ of continuous $1$-cocycles $\Phi:\Gamma_{F}\rightarrow\operatorname{Hom}(M,1+\mathfrak{m}_{A})$. Further, one may interpret $\widehat{Z}^{1}$ as $D^{\square}_{\overline{\rho}}$ for the group $G=\mathfrak{D}(M)\rtimes\underline{\Delta}$ and $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ given by $\overline{\rho}(\gamma)=(1,\pi(\gamma))$, where $\pi:\Gamma_{F}\rightarrow\Delta$ is the quotient map. The functor $\widehat{Z}^{1}$ is the analogue of $D^{\square}_{\mathbf{1}}$ and we have to find the ring pro-representing the functor $\widehat{Z}^{1}$.
@@
Our solution is motivated by Langlands’ paper [11] on his correspondence for tori, where he shows that the universal coefficient theorem induces an isomorphism
@@
corresponding to the fundamental class $[u_{E/F}]\in H^{2}(\Delta,E^{\times})$. Since every continuous representation $\Gamma_{F}\rightarrow G(A)$ with image in $(G/G^{0})(A)$ equal to $\Delta$ factors through the profinite completion of $W_{E/F}$, this motivates us to study the functor
@@ 1.2. Abstract setting
is represented by the group algebra $\mathbb{Z}[H_{1}(\Gamma_{1},M)]$.
@@
To prove 1.7 we verify that the action of $\Delta$ on $(\Gamma_{2}^{\mathrm{ab}}\oplus I_{\Delta})\otimes M$ induced by its action on the functor in (6) is isomorphic to $\mathcal{E}\otimes M$. This calculation is carried out in Section 4. The last part of Theorem 1.7 follows by observing that (1) remains an isomorphism if $\mathbb{C}^{\times}$ is replaced by any divisible group and for any $A$ there is a faithfully flat map $A\rightarrow B$ such that $B^{\times}$ is divisible.
@@
$\operatorname{Rep}^{\Gamma_{1}}_{G}\sslash G^{0}\overset{\cong}{\longrightarrow}\mathrm{PC}^{\Gamma_{1}}_{G}.$
@@
Let us assume that $\operatorname{Rep}^{\Gamma_{1}}_{G}(\mathcal{O})$ is non-empty, $G^{0}$ is split over $\mathcal{O}$ and $G/G^{0}$ is a constant group scheme. Let $\rho_{0}:\Gamma_{1}\rightarrow G(\mathcal{O})$ be a representation and let $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ be the subfunctor of $\operatorname{Rep}^{\Gamma_{1}}_{G}$ consisting of representations $\rho$ such that $\Pi\circ\rho=\Pi\circ\rho_{0}$, where $\Pi:G\rightarrow G/G^{0}$ is the quotient map. One may similarly define $\mathrm{PC}^{\Gamma_{1}}_{G,\pi}$. We show that the map
@@ 1.3. Producing a lift
Let us go back to the problem of exhibiting a lift of $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ to characteristic zero as claimed in 1.1. Using $\overline{\rho}$ instead of $\rho_{0}$, we may identify the restriction of $D^{\square}_{\overline{\rho}}$ to $\mathfrak{A}_{k}$ with the restriction of $\widehat{Z}^{1}$ to $\mathfrak{A}_{k}$. This allows us to conclude that
@@
$\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M\cdot([F:\mathbb{Q}_{p}]+1),$
@@ 1.5. Notation
If $X$ is a scheme then we denote the ring of its global sections by $\mathcal{O}(X)$. If $\Gamma$ is a group then we denote by $\mathcal{O}[\Gamma]$ its group algebra over a ring $\mathcal{O}$.
@@ 1.6. Acknowledgements
VP would like to thank Timo Richarz for a helpful discussion concerning the Langlands correspondence for tori. The authors would like to thank Toby Gee and James Newton for their comments on a draft version of this paper. The authors thank the anonymous referee for their helpful comments that improved the quality of the manuscript.
@@ 2. Generalised tori
It follows from [6, Proposition 3.1.3], that for a generalised torus $G$ the quotient $G/G^{0}$ is étale. Sean Cotner has shown in [7] that if $S$ is locally noetherian then for a smooth affine $S$-group scheme $G$ the identity component $G^{0}$ has reductive geometric fibers and $G/G^{0}$ is finite étale if and only if $G$ is geometrically reductive in the sense of [1, Definition 9.1.1]. In particular, this holds over $S=\operatorname{Spec}\mathcal{O}$.
@@
If $M$ is an abelian group then we denote by $\mathfrak{D}(M):=\operatorname{Spec}(\mathcal{O}[M])$ the diagonalisable group scheme, where the Hopf algebra structure on $\mathcal{O}[M]$ is given by addition and inversion in $M$. For any $\mathcal{O}$-algebra $A$, we have
@@ 3. Shapiro’s lemma
Let $\varphi:\Gamma_{2}\rightarrow V$ be a $1$-cocycle, and let $f_{\varphi}:\Gamma_{2}\rightarrow\mathcal{F}_{V}$ be the $1$-cocycle corresponding to $\varphi$, via the canonical isomorphism $\mathcal{F}_{V}\cong V$ and let ${\Phi_{\varphi}:\Gamma_{1}\rightarrow\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}{V}}$ be the function given by
@@
The map $\varphi\mapsto\Phi_{\varphi}$ depends on the choice of the coset representatives $\bar{c}$ and so it is not canonical. However once these representatives have been fixed it is immediate from the formulas that the map is functorial in $V$. If $\alpha:V\rightarrow W$ is a $\Gamma_{2}$-equivariant homomorphism of abelian groups equipped with $\Gamma_{2}$-action and $\psi\in Z^{1}(\Gamma_{2},W)$ is the image of $\varphi$ under the map $Z^{1}(\Gamma_{2},V)\rightarrow Z^{1}(\Gamma_{2},W)$ induced by $\alpha$ then $\Phi_{\varphi}$ maps to $\Phi_{\psi}\in Z^{1}(\Gamma_{1},\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}{W})$.
@@ 4. The space of $1$-cocycles
Let $V$ be an abelian group equipped with a trivial $\Gamma_{2}$-action. We may identify $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V$ with the space of functions $f:\Delta\rightarrow V$ where the action is given by $[g\cdot f](c)=f(cg)$ for all $g\in\Gamma_{1}$ and $c\in\Delta$. We also have an action of $\Delta$ on $\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}V$, which commutes with the action of $\Gamma_{1}$ and is given by $[d\cdot f](c)=f(d^{-1}c)$ for all $d\in\Delta$. This induces an action of $\Delta$ on $Z^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V)$, which is given explicitly by
@@
We want to understand the action of $\Delta$ on $B^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V)\oplus\operatorname{Hom}(\Gamma_{2},V)$ via this isomorphism. It follows from (8) that $[[d\ast\Phi](g)](dc)=[\Phi(g)](c)$. Hence
$[d\ast\Phi_{\varphi}](g)=\sum_{c\in\Delta}\varphi(\overline{c}g\overline{cg}^{-1})\mathbf{1}_{dc}.$
@@
$[[d\ast\Phi_{\varphi}](k)](1)=\sum_{c}\varphi(\overline{c}k\overline{c}^{-1})\mathbf{1}_{dc}(1)=\varphi(\overline{d^{-1}}k\overline{d^{-1}}^{-1})=\varphi(\overline{d}^{-1}k\overline{d})=:[d\cdot\varphi](k).$
We note that $\varphi((\overline{d}h)^{-1}k(\overline{d}h))=\varphi(\overline{d}^{-1}k\overline{d})$ for all $h,k\in\Gamma_{2}$, which justifies the third equality and we take the last equality as the definition $d\cdot\varphi$. Since $[[\Phi_{d\cdot\varphi}](k)](1)=[d\cdot\varphi](k)$ by 3.2, 3.1 implies that the cocycle $d\ast\Phi_{\varphi}-\Phi_{d\cdot\varphi}$ is a $1$-coboundary and thus there is a function $f:\Delta\rightarrow V$ depending on $d$ such that for all $g\in\Gamma_{1}$ we have
@@
Since $V$ is abelian we have $\operatorname{Hom}(\Gamma_{2},V)=\operatorname{Hom}(\Gamma_{2}^{\mathrm{ab}},V)$, where $\Gamma_{2}^{\mathrm{ab}}$ is the maximal abelian quotient of $\Gamma_{2}$. For $d,c\in\Delta$ we define $\kappa(d,c)$ to be the image of $\overline{d}\overline{c}\overline{dc}^{-1}$ in $\Gamma_{2}^{\mathrm{ab}}$. The map $\kappa:\Delta\times\Delta\rightarrow\Gamma_{2}^{\mathrm{ab}}$ is the $2$-cocycle corresponding to the extension
@@
see [4, Chapter XIV, §4, Remark]. The extension class of (17) gives an element in $\operatorname{Ext}^{1}_{\mathbb{Z}[\Delta]}(I_{\Delta},\Gamma_{2}^{\mathrm{ab}})$. One can show, that the image of this class in $H^{2}(\Delta,\Gamma_{2}^{\mathrm{ab}})$ is equal to the class of $\kappa$.
@@
where the action of $\Gamma_{1}$ on the last term is induced by its action on $\mathbb{Z}[\Delta]\otimes M$, which is given by $g\cdot(c\otimes m)=cg^{-1}\otimes m$, so that $[g\cdot f](x)=f(g^{-1}x)$ for all $x\in\mathbb{Z}[\Delta]\otimes M$. The diagonal action of $\Delta$ on $\mathbb{Z}[\Delta]\otimes M$ is given by $d\cdot(c\otimes m)=dc\otimes dm$, commutes with the action of $\Gamma_{1}$ and induces a (left) action of $\Delta$ on $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V)$, which commutes with the action of $\Gamma_{1}$. If $M=\mathbb{Z}$ with the trivial $\Delta$-action then the construction recovers the action of $\Delta$ on $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V$ considered earlier.
@@
As $\Gamma_{1}$-representations $(\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V))^{\Delta}$ is naturally isomorphic to $\operatorname{Hom}(M,V)$ with $\Gamma_{1}$ acting on it via $\Delta$.
@@
Let $b\in B^{1}(\Gamma_{1},\operatorname{Hom}(M,V))$ be a boundary with $b(g)=(g-1)\alpha$ for all $g\in\Gamma_{1}$. Its image in $B^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V))$ is the boundary $b^{\prime}$ given by $b^{\prime}(g)=(g-1)\vartheta(\alpha)$. The constant function $f^{\prime}(c)=\alpha$ for all $c\in\Delta$ is $\Gamma_{1}$-invariant thus
@@
If $V$ is any abelian group then for all $\Gamma_{1}$-modules $N$ and all $i\geq 0$ we have a canonical map
(21) $H^{i}(\Gamma_{1},\operatorname{Hom}(N,V))\rightarrow\operatorname{Hom}(H_{i}(\Gamma_{1},N),V),$
@@
Lemmas 4.6 and 4.8 give us the commutative diagram above. Moreover, if $V$ is divisible then by 4.7 the top row is exact, the last arrow in the top row is surjective and the first vertical arrow is an isomorphism. We deduce that for all divisible $V$ the maps $q:(\mathcal{E}\otimes M)_{\Delta}\twoheadrightarrow I_{\Delta}M$ and $\varphi_{\mathrm{nat}}:H_{1}(\Gamma,M)\rightarrow(\mathcal{E}\otimes M)_{\Delta}$ induce an exact sequence
@@
(25) $\mathbb{Z}[(\mathcal{E}\otimes M)_{\Delta}]\cong\mathbb{Z}[H_{1}(\Gamma_{1},M)][t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}],$
@@
The right hand side is an fpqc sheaf, since it is representable [15, Tag 023Q]. So it suffices to find for every ring $A$ and every homomorphism ${f:\mathbb{Z}[H_{1}(\Gamma_{1},M)]\to A}$ an fpqc cover of $A$ and a descent datum in the left hand side, which maps to a descent datum in the right hand side, which descends to $f$. By [17, Lemma 4.1.1] for every ring $A$ there is a faithfully flat map $A\to B$, such that $B^{\times}$ is divisible. By 4.7 the map